OpenStudy (anonymous):

Trigonometric Functions Solve the following equations for 0<=x<=360. cos(70-x)=0.5

OpenStudy (mathmale):

One way to begin would be to solve the equation cos X = 0.5. This has two solutions, one in the 1st quadrant and 1 in the 4th. Can you find the two X values? Once you have these X values, equate each to 70-x (one at a time) and solve for x. Be certain to check your answers (which are the x-values you've just found).

OpenStudy (anonymous):

The first value is 10 but i'm not sure how to find the second value

OpenStudy (mathmale):

what are y our two X values? X=60 (which produces x=10) is correct. What's your other X?

OpenStudy (mathmale):

Which angle X, in the fourth quadrant, has a cosine of 0.5?

OpenStudy (anonymous):

is it x=120?

OpenStudy (mathmale):

120 degrees is not in the fourth quadrant; it's in the second. Please try again. What angle X in the fourth quadrant has a cosine of 0.5?

OpenStudy (anonymous):

is it 360

OpenStudy (mathmale):

closer than before. However, Marc, cos 360 is 1. cox X is 0.5; what is x?

OpenStudy (anonymous):

x=60

OpenStudy (mathmale):

We need to concentrate on Q4. if you have a calculator, set it to degree mode and type in cos 300. what do you get?

OpenStudy (anonymous):

0.5

OpenStudy (mathmale):

right. so X in Q4 is 300. make sense?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

Thought I had it, but I've discovered a mistake in my own work. However, if we let 70-x = 60, x=10 as you said. Try this : let 70-x=300. What's (small) x?

OpenStudy (mathmale):

If small x is negative, we must discard it, since small x must be between 0 and 360.

OpenStudy (anonymous):

x=-230

OpenStudy (mathmale):

That's same as mine. Because it's negative, we feed it to the dogs.

OpenStudy (mathmale):

that leaves us with one final answer: small x = what?

OpenStudy (anonymous):

x=10

OpenStudy (mathmale):

right. very good.

OpenStudy (anonymous):

Thank you @mathmale :)

OpenStudy (mathmale):

My great pleasure! See you again.

OpenStudy (anonymous):

cos(70-x)=0.5 70-x=cos^{-1} 0.5 70-x= 60 x=10

OpenStudy (anonymous):

since cos, x is also in the 4th quadrant , thus x =360-10=350

OpenStudy (anonymous):

@mathmale it is supposed to be 130 (adding 360 to the angle wont change the value of cosine ) 360-230=130 thnx @mathmale